# General-relativistic magnetar magnetospheres in 3D with physics-informed neural networks

### v1

@@ 5.2 Outburst energetics
Expression (15) can be used to determine the available free energy in the magnetosphere, $E_{e}$ from equation (12), in terms of observables.
Figure 7 shows burst energetics versus maximum free energies ($\approx 0.3E^{(M)}_{0}$) using data collated in the catalogue of Coti Zelati et al. (2018) after correcting for the GR factors in (15). We use the Tolman-VII moment of inertia $I_{0}\approx 2MR^{2}(1-1.1M/R-0.6M^{2}/R^{2})^{-1}/7$ (Lattimer & Prakash, 2001) with a mass $M\approx 1.4M_{\odot}$ and radius $R=12$ km. Even with these reductions, we see that the maximum free energy is more than sufficient to fuel outbursts (below the dashed line depicting equality). One event, however, which is borderline comes from the ‘low-field’ magnetar SGR 0418+5729 (van der Horst et al., 2010): the release of $E_{\rm burst}\sim 2\times 10^{41}$ erg in the 2009 event (Rea et al., 2013) points towards multipolarity since the entire dipolar, free-energy budget of the magnetosphere may be too low if the inclination angle $\chi\sim\pi/2$ as inferred from the double-peaked light curves observed by the Rossi X-ray Timing Explorer (RXTE; Esposito et al., 2010). Moreover, local twist releases may be expected rather than global ones so that $E_{e}=0.3$ is overly generous. If only $\sim 3\%$ of the free energy was released instead (Fig. 5), this implies that $l>1$ multipoles would have to comprise $\sim 6$ times more energy than the dipole. For other cases, such as for the 2003 outburst from XTE J1810–197 ($E_{\rm burst}\approx 4\times 10^{42}$ erg), a dipolar free energy of $\lesssim 2\%$ would suffice.
Figure 7: Comparison between available dipolar free energies for the maximally stable configurations ($30\%$ of the total untwisted energy), as determined using the GR-revised relation (15) to infer the polar field strength via timing data for masses $M=(1.4\pm 0.2)M_{\odot}$, radii $R=12$ km, Tolman-VII moment of inertia, and inclinations $\chi\in[0,\pi/2]$, against burst energies from the catalogue of Coti Zelati et al. (2018). The dashed line depicts equality.
The most energetic flare recorded from a (Galactic) magnetar came from the December 2004 event from SGR 1806–20, with a total (isotropic) energy yield of up to $\sim 5\times 10^{46}$ erg depending on distance (Palmer et al., 2005; Terasawa et al., 2005). For an orthogonal rotator, the observed period and period derivative values yield a polar field estimate of $B_{p}\approx 7\times 10^{14}$ G, a factor $\sim 5$ smaller than the ‘canonical’ Newtonian, vacuum value ($B_{p}\approx 4\times 10^{15}$ G). The maximum free energy in the magnetosphere can be estimated as
@@
Thus, although smaller distances, beaming, smaller inclination angles, or a less compact star could be invoked to shift the observed energy below that of expression (18), multipolar components may be anticipated to explain the energetics.
@@ 5.3 Afterglow light curves
The story changes somewhat if using instead the GR expression for the spindown luminosity, which is augmented relative to the Newtonian case by a factor $\kappa=f_{R}^{2}/N_{R}^{4}$ where $f_{R}$ is the right-hand side of expression (13) and $N_{R}=e^{\nu(R)}$ (Rezzolla & Ahmedov, 2004). For $M/R=0.3$ we find $\kappa=22.4$, meaning that the inferred field should be considerably lower. Keeping the plateau duration fixed, Rowlinson et al. (2013) note that $B_{p}^{2}\propto L^{-1}$ in their fitting procedure and thus – all else being equal – the inferred field for the putative star born in GRB 070724A may reduce to $B_{p}\sim 6\times 10^{15}$ G or even lower with twists. While still large, such a value is easier to accept (for realistic energy conversion factors, the inferred $B_{p}$ decreases further; see Suvorov & Kokkotas, 2021).
@@ 6 Conclusions
In this paper, we have considered the problem of building general-relativistic and fully three-dimensional models of magnetar magnetospheres. Using a prescription for the twist profile (expression 9) based on the expectation that surface currents may develop in localised regions during the course of crustal evolution or otherwise, a sequence of models were considered of varying compactness and ‘hotspot’ particulars (e.g. Fig. 4). We find that the lowest-energy configurations (with respect to given boundary data) can house a free energy of at most $\sim 30\%$ regardless of the compactness (see Fig. 3), consistent with previous Newtonian studies (e.g. Akgün et al., 2018). The free energy excess decreases in more realistic 3D cases with localised twists to at most $\approx 5\%$. Regardless of the twist particulars we also find that that $E_{e}$ is hardly changed as a function of compactness (see Fig. 5). We argue in Sec. 5 that because the spindown luminosity scales strongly with GR effects (Rezzolla & Ahmedov, 2004), the energy budget available to fuel magnetar bursts may be considerably lower than previously estimated. For some events – notably the 2009 burst from SGR 0418+5729 (van der Horst et al., 2010) – the dipolar excess may in fact be insufficient even for a (likely unrealistic) maximally-twisted configuration (see Fig. 7). This provides further evidence for multipolar components in magnetar fields.
@@ Acknowledgements
We acknowledge support from the Prometeo excellence programme grant CIPROM/2022/13 funded by the Generalitat Valenciana. We also acknowledge support through the grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union, as well as the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1).
@@ Appendix A Details of the numerical implementation
The PINN solver is configured using the following hyperparameters. The network is divided into four subnetworks, each dedicated to predicting one of the four target functions $B_{r},B_{\theta},B_{\phi},$ and $\alpha$. For axisymmetric models, each subnetwork comprises two hidden layers with 20 neurons per layer, while for 3D models, each subnetwork uses three hidden layers with 40 neurons each. The training set consists of $N=2\times 10^{4}$ points for the axisymmetric case and $N=8\times 10^{4}$ points for the 3D case, with the points randomly sampled within $\mathcal{D}$ which extends from the stellar surface to radial infinity using compactified coordinates (see below).
@@
To ensure adequate coverage of the entire domain, the training points are resampled every 700 iterations. The training process begins with the Adam optimiser, used for the first 1000 iterations to approach the loss minimum. After this initial phase, we switch to the self-scaled Broyden method, which achieves significantly higher accuracy (often by several orders of magnitude) compared to commonly used optimization algorithms (Adam, BFGS, L-BFGS; see Urbán et al., 2025, for details). A typical training takes $\sim 10-20$ minutes on a NVIDIA H100 GPU with 80GB RAM or $\sim 40-60$ minutes on a NVIDIA GeForce GTX 1660 Ti GPU with 6GB RAM for each model. In the 3D case, this approach demonstrates faster performance (roughly by a factor 2) than the method used in Stefanou et al. (2023a) for $M=0$ cases, although a direct comparison is challenging due to differences in the underlying hardware (GPU versus CPU).
@@
In all cases, we aim at a loss function that reaches values of at most $10^{-6}$, with the vast majority of models reaching $\mathcal{L}\lesssim 10^{-8}$. The global error $\varepsilon$ in the PDE can be estimated as $\varepsilon\simeq\sqrt{\mathcal{L}}$ (Urbán et al., 2025), corresponding to an estimated precision of $\varepsilon\lesssim 10^{-4}$ as can be seen in Figure 8, which is deemed sufficient. Models with $\varepsilon\gtrsim 10^{-3}$ (to the right of dashed line in the figure) are discarded as they violate the FF condition (see Appendix B.1). It is worth highlighting that the global errors are typically smaller in cases with non-zero compactness for the same $\alpha_{*}^{\rm max}$, meaning that we expect GR effects to be correctly captured. The actual error in the solution is expected to be lower than the PDE error (again we direct the interested reader to Urbán et al., 2025).
@@ Appendix B Magnetic energies
×$\times\Big{[}2M(M-R)+R(2M-R)\log\left(1-\frac{2M}{R}\right)\Big{]},$
@@ B.1 Criteria to detect force-free breakdown
A useful diagnostic for whether a given numerical output is strictly FF comes from the virial theorem (Chandrasekhar, 1961). It implies that a FF field within a (semi-infinite) domain $V$, decaying sufficiently fast with radius [$B\lesssim\mathcal{O}(r^{-3/2})$ at infinity], satisfies
@@
Figure 9 shows the ratio of the relevant expressions, which are of order unity until $\alpha_{*}^{\rm max}\approx 2.5R^{-1}$. After this point some visible scatter appears meaning that, although the PINN solution is still optimal in some appropriate sense, it does not adequately satisfy the FF condition everywhere. It is for this reason we have considered models with $\alpha_{*}^{\rm max}\lesssim 2.5R^{-1}$ in the main text (the exact maximum depends on twist particulars and compactness). Curiously, a significant deviation from unity starts where the models in Stefanou et al. (2023a) stopped converging (for $M=0$) using an independent, Grad-Rubin method. For the axysimmetric models, the same logic was applied to select $\alpha_{*}^{\rm max}\lesssim 4R^{-1}$.

### v2

@@ 5.2 Outburst energetics
Figure 7: Comparison between dipolar magnetospheric energies (see Appendix B) and burst energies from the catalogue of Coti Zelati et al. (2018) for a number of sources (see overlaid legends). In cases of multiple bursts by a given object, only the brightest event is shown. The polar-dipole strengths are determined using the GR-revised relation (16) via the available timing data for masses $M=1.4M_{\odot}$, radii $R=12$ km, Tolman-VII moment of inertia, and inclinations spanning $0\leq\chi\leq\pi/2$: that corresponding to the ‘canonical’ value (17) is depicted by small rhomboids (see the top-left legend). The dotted, blue line depicts the maximum free energy available in cases with axisymmetric twists ($\approx 30\%$ of the total; Fig. 3), with the dashed, red line corresponding instead to cases with localised twists ($\approx 3\%$; Fig. 5).
Expression (16) can be used to determine the free energy in the magnetosphere in terms of observables. Recalling Fig. 6, we have that the total magnetospheric energy is lower by a factor $\sim 3$ relative to the timing-based Newtonian inference for typical masses and radii ($M/R\approx 0.17$). This prompts us to revisit magnetar burst energies, $E_{\rm burst}$, in a GR context.
Figure 7 shows $E_{\rm burst}$ values versus $E^{(0.17)}_{0}$ for several sources using data collated in the catalogue of Coti Zelati et al. (2018). The reported timing parameters ($P\dot{P}$) are used to estimate the GR-corrected polar-field strength using equation (16), from which we calculate the total energy (expression 31). Typical values of the energy are found by applying expression (17), as shown by the small rhomboids ($\chi=\pi/4)$, though we allow for a range of inclinations ($0\leq\chi\leq\pi/2$) which define horizontal ‘error bars’. In particular, orthogonal rotators spindown faster for a fixed dipole moment meaning that matching to a given $P$ and $\dot{P}$ implies a lower magnetospheric energy relative to inferences for $\chi=\pi/4$. We overlay lines in the Figure to show the maximum free energies for stable cases with axisymmetric ($E_{e}=0.3$; Fig. 3) or localised ($E_{e}=0.03$; Fig. 5) twists.
[…]
The most energetic flare recorded from a (Galactic) magnetar came from the December 2004 event from SGR 1806–20, with a total (isotropic) energy yield of up to $\sim 5\times 10^{46}$ erg depending on distance (Palmer et al., 2005; Terasawa et al., 2005). For an orthogonal rotator, the observed period and period derivative values yield a polar field estimate of $B_{p}\approx 7\times 10^{14}$ G, a factor $\sim 5$ smaller than the Newtonian, vacuum value ($B_{p}\approx 4\times 10^{15}$ G) from expression (18). The maximum free energy in the magnetosphere can be estimated as
@@
Thus, although smaller distances, beaming, smaller inclination angles, or a less compact star could be invoked to shift the observed energy below that of expression (19), multipolar components may be anticipated to explain the energetics. At least in the axisymmetric context, Akgün et al. (2018) explored twist profiles different to our expression (8) and established a free-energy upper limit of $\approx 25\%$ for stable cases, so that we expect $E_{e}=0.3$ to be a firm bound.
@@ 5.3 Afterglow light curves
The story changes somewhat if using instead the GR expression for the spindown luminosity (15). For $M/R=0.3$ we find $\kappa=22.4$, meaning that the inferred field should be considerably lower. In fact, this is a lower limit since rapidly rotating stars have even further reductions: taking a $\sim 10\%$ change in the $tt$-component of the metric as representative (see footnote 1), we anticipate $N_{R}$ to decrease by a further $\sim 5\%$ so that $\kappa\approx 28$ for a $P\sim 2$ ms object (see also figure 4 in Ruiz et al., 2014). Keeping the plateau duration fixed, Rowlinson et al. (2013) note that $B_{p}^{2}\propto L^{-1}$ in their fitting procedure and thus – all else being equal – the inferred field for the putative star born in GRB 070724A may reduce to $B_{p}\lesssim 6\times 10^{15}$ G or even lower with twists or rapid rotation. While still large, such a value is easier to accept (for realistic energy conversion factors, the inferred $B_{p}$ decreases further; see Suvorov & Kokkotas, 2021).
@@ 6 Conclusions
In this paper, we have considered the problem of building GR and fully three-dimensional models of magnetar magnetospheres. While GR can influence several aspects of magnetar phenomena, such as pair cascades and particle drag dynamics (Beskin, 1990; Philippov et al., 2015), our focus here is on the equilibrium of magnetic stresses. Using a prescription for the twist profile (expression 9) based on the expectation that surface currents may develop in localised regions during the course of crustal evolution or otherwise, a sequence of models were considered of varying compactness and ‘hotspot’ particulars (e.g. Fig. 4). We find that the stable configurations (with respect to given boundary data) can house a free energy of at most $\sim 30\%$ regardless of the compactness (see Fig. 3), consistent with previous Newtonian studies (e.g. Akgün et al., 2018). The free energy excess decreases in more realistic 3D cases with localised twists to at most $\approx 5\%$. Regardless of the twist particulars we also find that that $E_{e}$ is hardly changed as a function of compactness (see Fig. 5). We argue in Sec. 5 that because the spindown luminosity scales strongly with GR effects (Rezzolla & Ahmedov, 2004), the energy budget available to fuel magnetar bursts may be considerably lower than previously estimated. For some events – notably the 2009 burst from SGR 0418+5729 (van der Horst et al., 2010) – the dipolar excess may in fact be insufficient even for a (likely unrealistic) maximally-twisted configuration (see Fig. 7). This provides further evidence for multipolar components in magnetar fields.
@@ Acknowledgements
We thank José Carlos Olvera for valuable guidance on the RNS code and an anonymous referee for their constructive feedback, particularly for their suggestion that resulted in Appendix C. We acknowledge support from the Prometeo excellence programme grant CIPROM/2022/13 funded by the Generalitat Valenciana. We also acknowledge support through the grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union, as well as the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1).
@@ Appendix A Details of the numerical implementation
We describe here the hyperparameters used to configure the PINN solver. The network is divided into four subnetworks, each dedicated to predicting one of the four target functions $B_{r},B_{\theta},B_{\phi},$ and $\alpha$. For axisymmetric models, each subnetwork comprises two hidden layers with 20 neurons per layer, while for 3D models, each subnetwork comprises three hidden layers with 40 neurons each. The training set consists of $N=2\times 10^{4}$ points for the axisymmetric cases and $N=8\times 10^{4}$ points for the 3D cases, with the points randomly sampled within $\mathcal{D}$ which extends from the stellar surface to radial infinity in compactified coordinates (see below).
@@
To ensure adequate coverage in the entire domain, the training points are resampled every 700 iterations. The training process begins with the Adam optimiser, used for up to the first 1000 iterations to approach the loss minimum. After this initial phase, we switch to the self-scaled Broyden method, which achieves significantly higher accuracy (often by several orders of magnitude) compared to commonly used optimization algorithms (Adam, BFGS, L-BFGS; see Urbán et al., 2025, for details). We train the models for up to a maximum of 8000 total iterations A typical training takes $\sim 10-20$ minutes on a NVIDIA H100 GPU with 80GB RAM or $\sim 40-60$ minutes on a NVIDIA GeForce GTX 1660 Ti GPU with 6GB RAM for each model. In the 3D case, this approach demonstrates faster performance (roughly by a factor of 2) than the method used in Stefanou et al. (2023a) for Newtonian cases, although a direct comparison is challenging due to differences in the underlying hardware (GPU versus CPU).
In all cases, we aim at a loss function that reaches values of at most $10^{-6}$, with the majority of models reaching $\mathcal{L}\lesssim 10^{-8}$ (see Urbán et al., 2025, in particular their Appendix C, for a discussion about errors and accuracy in PINNs). The convergence plot of two relativistic 3D models is shown in Figure 8 as an illustrative example (orange and blue solid lines, same models as in the third column of Fig. 4). We see that the highly twisted model ($\alpha_{*}^{\rm max}=2.50R^{-1}$), which is at the limit of acceptable, stable solutions (see Appendix B.1), reaches a loss value $\sim 10^{-7}$ as opposed to the low twist model ($\alpha_{*}^{\rm max}=1.00R^{-1}$) which reaches $\sim 10^{-8}$. Similar convergence plots are obtained for each of the models presented in this work. Most of the models that are discarded due to violating the FF condition (see Appendix B.1 below) present high peaks and the loss reduction halts to a value higher than $10^{-6}$, reflecting the difficulty of the solver to find the correct solution. Such a case is represented by the green dashed line in the figure.
@@
The global error $\varepsilon$ in the PDE can be estimated as $\varepsilon\simeq\sqrt{\mathcal{L}}$ (Urbán et al., 2025), corresponding to an estimated precision of $\varepsilon\lesssim 10^{-4}$ as can be seen in Figure 9, which is deemed sufficient. Models with $\varepsilon\gtrsim 10^{-3}$ (to the right of dashed line in the figure) are discarded as they violate the FF condition (see Appendix B.1). It is worth highlighting that the global errors are typically smaller in cases with non-zero compactness for the same $\alpha_{*}^{\rm max}$. The actual error in the solution is expected to be lower than the PDE error (again we direct the interested reader to Urbán et al., 2025).
Figure 8: Convergence of the loss function with iterations for two relativistic 3D models with $M/R=0.25$. The periodic peaks correspond to resamplings of the training points. The green dashed line shows an example of a model with poor convergence, which is discarded.
@@ Appendix B Magnetic energies
$\times\Big{[}2M(M-R)+R(2M-R)\log\left(1-\frac{2M}{R}\right)\Big{]},$
@@ B.1 Criteria to detect force-free breakdown
A useful diagnostic for whether a given numerical output is strictly FF comes from the virial theorem (Chandrasekhar, 1961). It implies that a FF field within a (semi-infinite) domain $V$, decaying sufficiently fast with radius $B\lesssim\mathcal{O}(r^{-3/2})$ at infinity], satisfies
@@
Figure 10 shows the ratio of the relevant expressions, which are of order unity until $\alpha_{*}^{\rm max}\approx 2.5R^{-1}$. After this point some visible scatter appears meaning that, although the PINN solution is still optimal in some appropriate sense, it does not adequately satisfy the FF condition everywhere. It is for this reason we have considered models with $\alpha_{*}^{\rm max}\lesssim 2.5R^{-1}$ in the main text (the exact maximum depends on twist particulars and compactness). Curiously, a significant deviation from unity starts where the models in Stefanou et al. (2023a) stopped converging (for $M=0$) using an independent, Grad-Rubin method. For the axysymmetric models, the same logic was applied to select $\alpha_{*}^{\rm max}\lesssim 4R^{-1}$.
Appendix C Stability and safety factor
A critical factor in evaluating the stability of twisted magnetic fields is the so-called safety factor, which is defined as
$q_{\rm{sf}}=\frac{4\pi}{\alpha\ell},$ (33)
for a field line (or a thin bundle) with constant twist density $\alpha$ and length $\ell$ (Mahlmann et al., 2023; Rugg et al., 2024). The inverse of the safety factor approximately indicates how many times magnetic field lines spiral around a flux tube. When the safety factor falls to about one or lower, the twisted magnetic structure becomes prone to kink instabilities (Kruskal & Schwarzschild, 1954), potentially triggering magnetar flares or outbursts accompanied by rapid magnetic energy release. Thus, the safety factor can be used as a diagnostic tool to predict when a magnetospheric configuration becomes unstable.
Axisymmetric models are known to be degenerate with the high-energy branch being unstable due to the presence of detached magnetic islands (see section 4.1 and Figure 3). Indeed, as Figure 11 shows, the safety factor of the field lines forming the magnetic island (after $\alpha_{*}^{\rm max}\approx 1.8$) in the high energy solution is $q_{\rm{sf}}\ll 1$, confirming that this region is highly susceptible to magnetic instabilities. Models in the low energy branch have a safety factor well above unity everywhere in the magnetosphere, as expected for the stable solutions.
For 3D models, a similar argument can be used. Figure 12 shows the safety factor of a bundle of twisted field lines for two relativistic 3D models. The less twisted model (teal circles in the figure) maintains a safety factor well above one for all its field lines. Conversely, in the highly twisted model (orange triangles in the figure), a sub-bundle of field lines originating from the central region of the hotspot has $q_{\rm{sf}}\lesssim 1$, suggesting the possibility for instabilities to develop. Remarkably, field lines that have $q_{\rm{sf}}\lesssim 1$ appear in models which have $\alpha_{*}^{\rm{max}}\gtrsim 2.5R^{-1}$, the same value as the one that marks the breakdown of the FF condition according to Appendix B.1. This reinforces the notion that beyond a critical twist, no stable FF solutions exist.
Relativistic corrections may work in favour of stability, albeit slightly. Increasing compactness has the effect of contracting flux bundles in the magnetosphere (see Figure 4), which results in higher safety factors due to the decreased length of the field lines. Indeed, for a fixed $\alpha_{*}^{\rm{max}}$, we have observed that $q_{\rm{sf}}$ is higher for compact stars, and may not reach the $q_{\rm{sf}}\sim 1$ threshold in cases where the Newtonian counterpart does. A similar effect is illustrated in Fig. 10, where the deviation from equality of the two energy definitions is slightly lower in the highly compact case. A detailed analysis of the connection between compactness and stability is outside the scope of this study and is planned for future research.
Figure 11: Field line safety factor $q_{\rm{sf}}$ (expression 33) as a function of the twist density $\alpha$ for the degenerate axisymmetric solutions presented in Fig. 2 ($M/R=0.25$, $s=0.243$ and $m=3$, resulting in $\alpha_{*}^{\rm{max}}=1.86R^{-1}$). Only field lines with $\alpha>0.5R^{-1}$ and one of their footprints in the $\phi=\pi$ plane are shown for clarity. All field lines in the low-energy solution have $q_{\rm sf}\gg 1$, indicating stability. For the high-energy solution, a sudden jump in the safety factor to values $q_{\rm{sf}}\ll 1$ can be observed, which corresponds to the field lines that form the unstable magnetic island. Notice that for these lines $\alpha>\alpha_{*}^{\rm{max}}$.
Figure 12: Safety factor $q_{\rm{sf}}$ as a function of the twist density $\alpha$ for a bundle of magnetic field lines emerging from the hotspot in two different relativistic 3D models: $\alpha_{*}^{\rm{max}}=2R^{-1}$ (teal circles) and $\alpha_{*}^{\rm{max}}=3R^{-1}$ (orange triangles). Both models have compactness $M/R=0.25$. Only field lines with $\alpha>0.5R^{-1}$ are shown for clarity.